OpenStudy (anonymous):

Sue deposited $3000 in a savings account earning 2% interest over 8 years. What was the total interest Sue earned?

OpenStudy (solomonzelman):

simple interest, or compound interest?

OpenStudy (anonymous):

simple

OpenStudy (solomonzelman):

Ok, then you will have to find the interest earned for the first year, then multiply that times 8 (for a total 8 year interest).

OpenStudy (solomonzelman):

your interest is 2% of 3000, that is == \(\large\color{black}{ \displaystyle 3000 \times 0.02 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3000 \times 0.02 =? }\) (you tell me)

OpenStudy (anonymous):

60.

OpenStudy (anonymous):

sorry it is 60

OpenStudy (anonymous):

@SolomonZelman